# Collective action and entanglement of magnetically active liquid crystal elastomer ribbons

### v1

@@
Abstract
@@
Interactions between active individuals in animal collectives lead to emergent responses that remain elusive in synthetic soft matter. Here, shape-morphing polymers are used to create bio-inspired transient solids that self-assemble with controlled mechanical properties and disassemble on demand. Dilute-suspensions of magnetic, heat-responsive liquid crystal elastomer ribbons mechanically interlock, inducing reversible aggregation. A mathematical model is developed that sheds light on the role of topological mechanisms in aggregation. Aggregation was favored for ribbons with moderate curvature at C above crosslinking temperature as compared to flat ribbons or higher curvature ribbons at higher temperatures. The ribbon suspensions reversibly transition between fluid- and solid-like states, exhibiting up to 6 orders-of-magnitude increase in the storage moduli of the entangled aggregates compared with the liquid dispersions. Controlled dissociation is induced by imparting kinetic energy to the individual ribbons at high magnetic field rotation speeds ( RPM). Ribbon shape and the medium in which dissociation occurs were shown to govern disassembly. Imparting dynamic collective behaviors into synthetic systems may enable a range of potential applications from bio-inspired soft robotics to injectable biomaterials.
@@
Keywords liquid crystal elastomers collective action active matter entangled matter actuators stimuli-responsive polymers soft robotics
1 Introduction
@@
Materials capable of self-assembly and disassembly in response to external cues hold promise to enable new functionality across engineering disciplines [1, 2, 3]. Such autonomous behaviors may offer novelty in fabrication and processing practices of bulk functional material, as well as in soft robotics, in which a swarming behavior of individual units can be manipulated to result in a stable solid structure [3, 4]. If such structures do not posses long range order (amorphous), they may further hold the benefit of requiring low levels of control over the assembly process to produce a multitude of forms with consistent properties [5, 6].
@@
One interesting approach towards building materials that self-assemble and disassemble is to rely on interactions between units that are able to move independently. The general term used for such systems is active matter. Such systems are composed of individual active units (across a variety of scales) that are far from equilibrium, which consume energy to generate motion or mechanical work [7, 8]. In many cases, the propelling of individual elements drives the self-assembly of the units into a structure or aggregate leading to collective behavior and emergent properties. If the cohesive forces between the aggregate elements are a result of the continuous supply of energy to the elements, the structure will break down upon consuming the source or the cutoff of the energy supply. In some cases, collisions may result in jamming or entanglement which can lead to metastable assembly due to the frustrated interaction between the units [9, 8].
@@
Some interesting sources of inspiration to cohesive interactions between active individuals that lead to collective mechanical behaviors can be found in the animal kingdom [8]. Notable examples are fire ants (Solenopsis invicta) [10], honeybees (Apis mellifera) [11], plant–animal worms (Symsagittifera roscoffensis) [12], and California blackworms (Lumbriculus variegatus) [13]. These species mechanically interlock (in response to environmental conditions) into cohesive masses that manifest various collective behaviors such as buoyant raft formation [14], colony unification [15] and thermal regulation [16]. By nature, these masses are porous structures with viscoelastic mechanical behaviors [17] that exhibit adaptive self-assembly and reorganization. These properties stem from each member’s ability to move within the confines of the structure, or in-and-out by physically detaching and reattaching to its neighbours [15, 18, 19, 20]. Moreover, their ability to reversibly transition between fluid- and solid-like states allows them to be uniquely adaptable and, as a mechanical system, to operate with a large number of degrees of freedom [21, 22]. In the context of a robotic system or a synthetic functional device made with a clear engineering application in mind, such capabilities have only been partially leveraged.
@@
To-date, several synthetic realizations of interlocked aggregates have been realized. For example, a variety of granular and swarming robotic systems that self-assemble into two dimensional aggregates using loose magnetic coupling have been proposed. These systems were able to demonstrate transitions between collective states with solid- and liquid-like properties [4] and exhibited emergent collective properties such as locomotion and transport [23, 24]. However, this approach uses bulky robotic components, each requiring relatively complex integrated electronic circuits and controls. Other approaches employ the unique attributes of responsive polymeric materials. The use of polymeric materials can greatly enhance the simplicity of the responsive unit. Abdelrahman et al. [6] demonstrated the creation of transient viscoelastic aggregates from a range of functional polymer ribbons. Such materials included liquid crystal elastomers (LCE) which are loosely crosslinked responsive polymer networks that reversibly change their shape in response to external stimuli such as heat or light. However, the aggregation required a temperature transient and large concentrations of overlapping entangling units for aggregate formation.
@@
In this work, we describe an approach to create transient, responsive solids from dilute suspension of mobile shape-changing polymeric ribbons. Specifically, we utilize the large-scale motion (i.e., much greater than a single body length) exhibited by magnetic and shape changing LCE ribbons in response to a rotating uniform magnetic field. When heated, ribbons change shape with prescribed amounts of bending and twisting, Figure 1a. Then, the rotation of the magnetic field induces aggregation through the entanglement of individual ribbons together into clusters, e.g., Figure 1b and Supplementary Video section A.1. We describe the thermal and magnetic response of individual ribbons, and track the aggregation kinetics of ribbon dispersions. Following the rotation of the magnetic field at elevated temperatures (65 ∘C, Figure 1c), the aggregates have a greater storage modulus than loss modulus up to shear strains of approximately 3%, constituting a solid-like behaviour, Figure 1d. The viscoelastic mechanical and rheological properties of the formed aggregates are found to vary with the single unit shape. Finally, we examine the critical conditions required for controlled aggregate disassembly.
2 Results and discussion
2.1 Thermal and Magnetic response
@@
LCE ribbons that respond to heat by changing shape and respond by moving in an external magnetic field were fabricated. The response to the magnetic field is conferred via a composite LCE domain at the edge of the ribbon filled with magnetic particles, Figure 1a and Supplementary Figure S1. The remainder of the LCE ribbon is shape responsive. The nematic director rotates by 90 ∘ through the film thickness (twisted nematic, Figure 2a), and heating causes the ribbon to morph from flat to 3D shapes with controlled degrees of bend and twist [25, 26]. Grazing incidence WAXS measurement on monodomain nematic film samples (uniform through the thickness) validated a high degree of molecular alignment along the prescribed orientation (Figure S1). The thermally induced transition from the nematic phase towards the isotropic phase locally induces a contraction along the nematic director that is accompanied by an expansion in the perpendicular direction. The amount of contraction is prescribed by the change in temperature. The most shape change occurs when heating above the nematic to isotropic transition temperature, . Twisted nematic ribbons with the nematic director at each face offset by an angle to the principal axes of the ribbon, Figure 2b, exhibit controlled bending and twisting [25, 26]. The evolution of curvature with temperature for ribbons cut at different offset angles is shown in Figure 2d, and the amount of twist (around the long axis of the ribbon) per unit length at different temperature is shown in Figure 2e. Ribbons that were not cut at an offset angle, , only bend, show an approximately linear rate of increase in curvature of approximately 0.014 mm-1/∘C. As expected, increasing the offset angle shows a decrease in curvature and an increase in the amount of observed twist. Mild offset angles of showed similar curvatures to with an increased amount of twist. The twisting of the ribbon gives rise to a helical shape, Figure 2b, with the pitch proportional to the amount of twist. Specifically, for a 16 mm ribbon at 65 ∘C, there is a quarter twist along the long axis of the ribbon. As temperature increases towards , the curvature of the ribbon increases, Figure 1a. Thus, the coil becomes tighter with decreases in pitch, length, and diameter.
Once the ribbon has acquired its actuated shape under the prescribed temperature, rotation of an external magnetic field is applied to induce the motion of the ribbon. The motion of the ribbon is a result of multiple factors including the magnitude of the magnetization in the ribbon, the external magnetic field, the geometry of the ribbon, and the viscosity of the media.
@@
Magnetic actuation of different ribbon shapes, caused by heating to different temperatures, induces trajectories that vary in space and time (Supplementary Figure S2 and Supplementary Video section A.1). To obtain a measure of the response of ribbons to an applied rotating field, the position of the magnetic domain was tracked in time in the plane (parallel to the fluid surface). The collected trajectories were filtered to a small range around the magnetic field rotation frequency, i.e., approximately 1.2 Hz or 72 RPM.
@@
The measured actuation paths of the ribbons consist of repeatable patterns that indicate three distinct motion regimes. The first pattern is characterized by many acute-angled features, 40 ∘C in Figures 2f-h. In this regime the ribbon frequently changes its direction of motion. The magnetic domain proceeds in a side-to-side flapping-like motion (e.g., Supplementary Video section A.1). This pattern of motion typically occurs when the ribbon is flat, and the magnetic moment is unable to make the ribbon bend or rotate about its axis before the external field completes a full rotation. The second motion regime is more circular, 65 ∘C-120 ∘C in Figures 2g-i, indicating that on the time scale related to the magnetic field rotation, the ribbon is able to follow the field without changing directions (e.g., Supplementary Video section A.1). The third motion regime contains more relaxed angular features oriented in different directions 65 ∘C-120 ∘C in Figure 2f. This pattern indicates a transition regime in which the ribbon is able to partially follow the field rotation but still displays a sizeable, albeit smaller, amount of direction changes per rotation cycle.
@@
2.2 Assembly and aggregation kinetics
Magnetic actuation of many curved ribbons in a fluid induces isothermal aggregate assembly from dispersed states. In each of the following experiments, forty ribbons were initially placed in a dish filled with viscous silicone oil, corresponding to a LCE packing density of 0.002 mg/mm3. The ribbons were placed in the oil such that there was no overlap between them (, Figure 3a). The motion induced by the magnetic field at low rotation speeds (72 RPM) facilitated entanglement of pairs of individual ribbons. Pairs subsequently grew into clusters through the annexation of additional ribbons or ribbon clusters, Supplementary Video section A.1.
Aggregation varies based on unit shape which is prescribed by the temperature. Time lapse images of dispersions of ribbons cut at () at different temperatures show different aggregation behavior, Figure 3a. At all temperatures that induce shape change, aggregation is observed. At a common actuation temperature of 65∘C, ribbons that bend and/or twist aggregate while those that remain flat (polydomain) do not. The time evolution of the # of clusters for offset angle at different temperatures (C) is presented in Figure 3b. At 40 ∘C (red line), the ribbons remained flat and exhibited discontinuous flapping motion, similar to the motion pattern exhibited by single ribbons in Figure 2h. The lack of bending means no aggregation is observed (Figure 3a and Supplementary Video section A.1).
@@ 2.2 Assembly and aggregation kinetics
A low number of resulting clusters provides an indication of ‘successful’ entanglement. However, to determine which of the above scenarios occurred, it is important to account for the distribution of cluster sizes. To account for this measure, we can consider the average number of ribbons in each cluster. The weighted average cluster size in the experiments presented in Figure 3b (C) is presented in Figure 3c. The C ribbon system exhibited a late-time value of clusters which had an average cluster size of ribbons/cluster, this constitutes a yield of merely of the initial number of ribbons. The C exhibited a late-time value of clusters which had an average cluster size of ribbons/cluster, constituting a yield of the initial number of ribbons. This is also evident from time lapse images presented at the rightmost column ( s) of Figure 3a. Thus, a cluster size approaching the initial number of units will indicate a successful entanglement that has resulted in the formation of a single mass.
@@
The shape of ribbons at a common temperature resulting from varying the offset angle is also expected to affect aggregate formation. We consider three different actuated shapes based on varying the offset angle, , exhibiting different amounts of curvature and twist. As a control, we consider a polydomain ribbon, which does not exhibit any actuation (i.e., remains flat) at elevated temperature. Time lapse images of dispersions of ribbons with different shapes at 65∘C are shown in Figure 3d. The time evolution of the # of clusters for different offset angles at 65∘C is presented in Figure 3e, and the average number of ribbons per cluster is presented in Figure 3f. Again, the flapping motion performed by the polydomain ribbons (red line) did not result in any bending from the ribbons (Supplementary Video section A.1) which did not result in entanglement.
@@
The kinetics of aggregation of different ribbon systems that aggregate did not exhibit any discernible differences. Although some variation is apparent, no substantial difference in kinetics or average cluster size was observed between the different angles at 65∘C. In all systems, two aggregation rates were observed, a fast drop in # of clusters lasting approximately 20 s followed by a slower decrease lasting approximately 80 s. In all systems, the formed aggregates occupied a volume approximately 80-90% smaller than in their dispersed state, Figure S3. This indicates that a continuous motion in circular patterns does result in successful entanglement. Similar observations were further reported by Patil et al. [13] in a computational study of active filaments representing dynamic reversible entanglement in the California blackworms. The evident considerable entanglement achieved in all systems in which the two of our preliminary conditions are fulfilled, i.e., continuous synchronicity with the field rotation and moderate curvature, can be considered to support our preliminary hypothesis.
2.3 Mathematical model for dynamic entanglement
@@
By developing a mathematical model for the ribbon dynamics we can identify the governing mechanisms underlying the aggregation process. In particular, this model allows us to understand the role of topological and non-topological mechanisms in aggregation. Our model treats the ribbons as rigid filaments moving in a cylindrical domain with diameter and height , where is the length of a ribbon. To promote entanglement, the process is started from a random initial condition in which the center of mass of each ribbon is in the same plane. Each ribbon then experiences constant driving torques, analogous to the forces on the magnetic domain in the LCE ribbons, along with an external stochastic drift force. The ribbons also feel contact forces and torques arising from their interactions [27, 28] (Supplementary Information, Supplementary Video section A.1). These contact forces typically include an adhesion term which prevents touching ribbons from separating unless their separation force exceeds a critical value. To understand how topological and non-topological parameters govern the aggregation process, we vary filament shape (Figure 4a; top row and bottom row) and adhesion strength (Figure 4a-c and Figure 4d-f).
@@
Filaments of different shapes exhibit different propensities to entangle with other, and thus help illustrate how topology affects aggregation. Here, two filament shapes are considered: a low curvature shape, corresponding to ribbons at C (Figure 4a, top row), and a high curvature shape, corresponding to ribbons at C (Figure 4a, bottom row). When adhesion forces are switched on, and the filaments are ‘sticky’, these different shapes give rise to strikingly different aggregation dynamics (Figure 4b) which resemble the experimental results in Figure 3b.
@@
To understand the topological origin of this difference, we introduce the excess linking per filament (Figure 4c), which quantifies the entanglements in a multi-filament system. This quantity is built out of the open linking number, , which captures pairwise entanglement between the ’th and ’th filament [13, 29]. For filaments, the excess linking per filament is the average amount of linking with a single filament, relative to the average amount of linking found in a random configuration of filaments (Supplementary Information). At , the low curvature filaments exhibit a much more robust amount of entanglement than the high curvature filaments, as expected from their larger cluster sizes (Figure 4b). Previous work has indicated that an open linking number with magnitude approximately matches an intuitive notion of entanglement between two filaments [13]. Since the high curvature filaments achieve an excess linking per filament of approximately (Figure 4c), this suggests that these filaments effectively pair off into clusters of size 2. This is consistent with the cluster number results (Figure 4b), which show the high curvature filaments forming clusters from initial filaments.

### v2

# Supp_matter Collective action and entanglement of magnetically active liquid crystal elastomer ribbons SUMMARY

@@
Supp_matter
Collective action and entanglement of magnetically active liquid crystal elastomer ribbons
SUMMARY
@@
Interactions between active individuals in animal collectives lead to emergent responses that remain elusive in synthetic soft matter. Here, shape-morphing polymers serve to create bio-inspired transient solids that self-assemble with controlled mechanical properties and disassemble on demand. Dilute-suspensions of magnetic, heat-responsive liquid crystal elastomer ribbons mechanically interlock, inducing reversible aggregation. A mathematical model is developed that sheds light on the role of topological mechanisms in aggregation. Aggregation was favored for ribbons with moderate curvature as compared to flat ribbons or higher curvature ribbons. The ribbon suspensions reversibly transition between fluid- and solid-like states, exhibiting up to 6 orders-of-magnitude increase in the storage moduli of the entangled aggregates compared with the liquid dispersions. Controlled dissociation is induced by imparting kinetic energy to the individual ribbons at high magnetic field rotation speeds ( Hz). Imparting dynamic collective behaviors into synthetic systems may enable potential applications from bio-inspired soft robotics to injectable biomaterials.
KEYWORDS
@@
liquid crystal elastomers, active matter, collective action, entangled matter, actuators, stimuli-responsive polymers, soft robotics
INTRODUCTION
@@
Materials capable of self-assembly and disassembly in response to external cues hold promise to enable new functionality across engineering disciplines 1, 2, 3. Such autonomous behaviors may offer novelty in fabrication and processing practices of bulk functional material, as well as in soft robotics, in which a swarming behavior of individual units can be manipulated to result in a stable solid structure 3, 4. If such structures do not posses long range order (amorphous), they may further hold the benefit of requiring low levels of control over the assembly process to produce a multitude of forms with consistent properties 5, 6.
@@
One interesting approach towards building materials that self-assemble and disassemble is to rely on interactions between units that are able to move independently. The general term used for such systems is active matter. Such systems are composed of individual active units (across a variety of scales) that are far from equilibrium, which consume energy to generate motion or mechanical work 7, 8. In many cases, the propelling of individual elements drives the self-assembly of the units into a structure or aggregate leading to collective behavior and emergent properties. If the cohesive forces between the aggregate elements are a result of the continuous supply of energy to the elements, the structure will break down upon consuming the source or the cutoff of the energy supply. In some cases, collisions may result in jamming or entanglement which can lead to metastable assembly due to the frustrated interaction between the units 9, 8.
@@
Some interesting sources of inspiration to cohesive interactions between active individuals that lead to collective mechanical behaviors can be found in the animal kingdom 8. Notable examples are fire ants (Solenopsis invicta) 10, honeybees (Apis mellifera) 11, plant–animal worms (Symsagittifera roscoffensis) 12, and California blackworms (Lumbriculus variegatus) 13. These species mechanically interlock (in response to environmental conditions) into cohesive masses that manifest various collective behaviors such as buoyant raft formation 14, colony unification 15 and thermal regulation 16. By nature, these masses are porous structures with viscoelastic mechanical behaviors 17 that exhibit adaptive self-assembly and reorganization. These properties stem from each member’s ability to move within the confines of the structure, or in-and-out by physically detaching and reattaching to its neighbours 15, 18, 19, 20. Moreover, their ability to reversibly transition between fluid- and solid-like states allows them to be uniquely adaptable and, as a mechanical system, to operate with a large number of degrees of freedom 21, 22. In the context of a robotic system or a synthetic functional device made with a clear engineering application in mind, such capabilities have only been partially leveraged.
@@
To-date, several synthetic realizations of interlocked aggregates have been realized. For example, a variety of granular and swarming robotic systems that self-assemble into two dimensional aggregates using loose magnetic coupling have been proposed. These systems were able to demonstrate transitions between collective states with solid- and liquid-like properties 4 and exhibited emergent collective properties such as locomotion and transport 23, 24. However, this approach uses bulky robotic components, each requiring relatively complex integrated electronic circuits and controls. Other approaches employ the unique attributes of responsive polymeric materials. The use of polymeric materials can greatly enhance the simplicity of the responsive unit. Wie et al., 25, 26 demonstrated collective action of magnetic assemblies of anisotropic composite polymer robots. They have shown that the exertion of kinetic energy through controlled rotation of magnetic fields induces swarming behaviors that can operate in both dry and wet environments. These swarms further demonstrated versatile task execution such as controlled assembly, disassembly, swimming, stirring and object manipulation. Abdelrahman et al. 6 demonstrated the creation of transient viscoelastic aggregates from a range of functional polymer ribbons. Such materials included liquid crystal elastomers (LCE) which are loosely crosslinked responsive polymer networks that reversibly change their shape in response to external stimuli such as heat or light. However, the aggregation required a temperature transient and large concentrations of overlapping entangling units for aggregate formation. A combination of the latter concepts utilizing functional polymers can offer a versatile platform for the creation of multifunctional programmable materials that exhibit a wide array of emergent collective behaviors.
@@
In this work, we describe an approach to create transient, responsive solids from dilute suspension of mobile shape-changing polymeric ribbons. Specifically, we utilize the large-scale motion (i.e., much greater than a single body length) exhibited by magnetic and shape changing LCE ribbons in response to a rotating uniform magnetic field. When heated, ribbons change shape with prescribed amounts of bending and twisting, Figure 1a. Then, the rotation of the magnetic field induces aggregation through the entanglement of individual ribbons together into clusters, e.g., Figure 1b and Supplementary Video List of supporting videos. We describe the thermal and magnetic response of individual ribbons, and track the aggregation kinetics of ribbon dispersions. Following the rotation of the magnetic field at elevated temperatures (65 ∘C, Figure 1c), the aggregates have a greater storage modulus than loss modulus up to shear strains of approximately 3%, constituting a solid-like behaviour, Figure 1d. The viscoelastic mechanical and rheological properties of the formed aggregates are found to vary with the single unit shape. Finally, we examine the critical conditions required for controlled aggregate disassembly.
RESULTS
Thermal and Magnetic response
@@ RESULTS
LCE ribbons that respond to heat by changing shape and respond by moving in an external magnetic field were fabricated. The response to the magnetic field is conferred via a composite LCE domain at the edge of the ribbon filled with magnetic particles, Figure 1a and Supplementary Figure S1. The remainder of the LCE ribbon is shape responsive. The nematic director rotates by 90 ∘ through the film thickness (twisted nematic, Figure 2a), and heating causes the ribbon to morph from flat to 3D shapes with controlled degrees of bend and twist 27, 28. Grazing incidence WAXS measurement on monodomain nematic film samples (uniform through the thickness) validated a high degree of molecular alignment along the prescribed orientation (Figure S1). The thermally induced transition from the nematic phase towards the isotropic phase locally induces a contraction along the nematic director that is accompanied by an expansion in the perpendicular direction. The amount of contraction is prescribed by the change in temperature. The most shape change occurs when heating through the nematic to isotropic transition temperature, noted as . In these materials some order is always retained on heating as such this transition is sometimes called the nematic-paranematic transition29. Twisted nematic ribbons with the nematic director at each face offset by an angle to the principal axes of the ribbon, Figure 2b, exhibit controlled bending and twisting 27, 28. The evolution of curvature with temperature for ribbons cut at different offset angles is shown in Figure 2d, and the amount of twist (around the long axis of the ribbon) per unit length at different temperature is shown in Figure 2e. Ribbons that were not cut at an offset angle, , only bend, show an approximately linear rate of increase in curvature of approximately 0.014 mm-1/∘C. As expected, increasing the offset angle shows a decrease in curvature and an increase in the amount of observed twist. Mild offset angles of showed similar curvatures to with an increased amount of twist. The twisting of the ribbon gives rise to a helical shape, Figure 2b, with the pitch proportional to the amount of twist. Specifically, for a 16 mm ribbon at 65 ∘C, there is a quarter twist along the long axis of the ribbon. Close-up images of actuated ribbons at different temperatures and offset angles are presented in Supplementary Figure S2. As temperature increases towards , the curvature of the ribbon increases, Figure 1a. Thus, the coil becomes tighter with decreases in pitch, length, and diameter.
Once the ribbon has acquired its actuated shape under the prescribed temperature, rotation of an external magnetic field is applied to induce the motion of the ribbon. The motion of the ribbon is a result of multiple factors including the magnitude of the magnetization in the ribbon head ( emu/g, Supplementary Figure S1), the external magnetic field ( mT, Supplementary Figure S3), the geometry of the ribbon, and the viscosity of the media. The ratio of inertial force to viscous force acting on the rotating swimming ribbon in a viscous fluid can be estimated by the rotational form of the Reynolds number 30, . Substitution of the parameters of the actuated ribbon at 65 ∘C and the viscous fluid yields an approximate value of , indicating a close balance between the forces and a laminar flow regime. Additional heating of the system will result in an increase in curvature (squared decrease in radius) and a proportional reduction in viscosity, indicating that this provides an upper bound for the relative magnitude of inertial forces. To evaluate the effect of rotation on the ribbon shape, we measured the change in curvature of individual ribbons in the presence of the rotating magnetic field. Ribbons exhibited an average % reduction in curvature during rotation at 1.2 Hz compared to the rest state ().
@@
Magnetic actuation of different ribbon shapes, caused by heating to different temperatures, induces trajectories that vary in space and time (Supplementary Figure S4 and Supplementary Video List of supporting videos). To obtain a measure of the response of ribbons to an applied rotating field, the position of the magnetic domain was tracked in time in the plane (parallel to the fluid surface). The collected trajectories were filtered to a small range around the magnetic field rotation frequency, i.e., approximately 1.2 Hz (72 RPM).
@@
The measured actuation paths of the ribbons consist of repeatable patterns that indicate three distinct motion regimes. The first pattern is characterized by many acute-angled features, 40 ∘C in Figures 2f-h. In this regime the ribbon frequently changes its direction of motion. The magnetic domain proceeds in a side-to-side flapping-like motion (e.g., Supplementary Video List of supporting videos). This pattern of motion typically occurs when the ribbon is flat, and the magnetic moment is unable to make the ribbon bend or rotate about its axis before the external field completes a full rotation. The second motion regime is more circular, 65 ∘C-120 ∘C in Figures 2g-i, indicating that on the time scale related to the magnetic field rotation, the ribbon is able to follow the field without changing directions (e.g., Supplementary Video List of supporting videos). The third motion regime contains more relaxed angular features oriented in different directions 65 ∘C-120 ∘C in Figure 2f. This pattern indicates a transition regime in which the ribbon is able to partially follow the field rotation but still displays a sizeable, albeit smaller, amount of direction changes per rotation cycle.
@@
Assembly and aggregation kinetics
Magnetic actuation of many curved ribbons in a fluid induces isothermal aggregate assembly from dispersed states. In each of the following experiments, forty ribbons were initially placed in a dish filled with viscous silicone oil, corresponding to a LCE packing density of 0.002 mg/mm3 (or a volume fraction of 0.0011 mm3/mm3). The ribbons were placed in the oil such that there was no overlap between them (, Figure 3a). The oil submersion allows for maintaining a constant temperature throughout the system, but it may also act to reduce adhesion between the ribbons, and supports motion in 3D by reducing effects of gravity. The motion induced by the magnetic field at low rotation speeds (1.2 Hz or 72 RPM) facilitated entanglement of pairs of individual ribbons. Pairs subsequently grew into clusters through the annexation of additional ribbons or ribbon clusters, Supplementary Video List of supporting videos. The side view in Supplementary Video List of supporting videos shows that the process occurs on the bottom of the dish in a fully submersed state, such that the surface tension of the oil does not affect the process.
Entanglement results from curved ribbons threading into each other. The rotation of the ribbons in the dish is accompanied by collisions of rotating individual ribbons or ribbon clusters. Such collisions result in topological interaction in which the active magnetic domain of one ribbon threads into the curved body of another and they begin to rotate together until interaction repeats with an additional ribbon and the cluster grows, Supplementary Video List of supporting videos and Supplementary Video List of supporting videos. Although interaction of the magnetic domains can occur in specific orientations, the weak forces between the magnetic domains (Supplementary Figure S1) are not sufficiently strong to hold the structure together or withstand the forced motions induced by the rotating field in the absence of ribbon curvature. This is further true for the cohesive interaction between the non-magnetic ribbon bodies, supporting the assumption that capillarity (e.g., Ref 31) does not play a dominant role in interaction. To illustrate this, the entangled structure can be cooled so that the ribbons return to a flat shape, and the magnetic field rotation is subsequently turned on again. Despite the close proximity of the magnetic domains and the ribbon bodies to each other, cooling and restarting the magnetic field results in a redispersion of the ribbons along the dish propelled by the previously described flapping motion (Supplementary Video List of supporting videos and Supplementary Video List of supporting videos). These findings lead us to hypothesize that the main contribution for cohesive interaction stems from static friction between ribbon surfaces. Additionally, we observe that ribbon clusters tend to rotate as a somewhat rigid body with minimal relative motion of ribbons within the cluster.
Aggregation varies based on unit shape which is prescribed by the temperature. Time lapse images of dispersions of ribbons cut at () at different temperatures show different aggregation behavior, Figure 3a. At all temperatures that induce shape change, aggregation is observed. At a common actuation temperature of 65∘C, ribbons that bend and/or twist aggregate while those that remain flat (polydomain) do not. The time evolution of the # of clusters for offset angle at different temperatures (C) is presented in Figure 3b. At 40 ∘C (red line), the ribbons remained flat and exhibited discontinuous flapping motion, similar to the motion pattern exhibited by single ribbons in Figure 2h. The lack of bending means no aggregation is observed (Figure 3a and Supplementary Video List of supporting videos).
@@
A low number of resulting clusters provides an indication of ‘successful’ entanglement. However, to determine which of the above scenarios occurred, it is important to account for the distribution of cluster sizes. To account for this measure, we can consider the average number of ribbons in each cluster. The weighted average cluster size in the experiments presented in Figure 3b (C) is presented in Figure 3c. The C ribbon system exhibited a late-time value of clusters which had an average cluster size of ribbons/cluster, this constitutes a yield of merely of the initial number of ribbons. The C exhibited a late-time value of clusters which had an average cluster size of ribbons/cluster, constituting a yield of the initial number of ribbons. This is also evident from time lapse images presented at the rightmost column ( s) of Figure 3a. Close-up images of aggregates entangled at different temperatures, C, are presented in Supplementary Figure S6a. Thus, a cluster size approaching the initial number of units will indicate a successful entanglement that has resulted in the formation of a single mass. The field rotation speed was shown to affect aggregation quality with rotation speeds below 1.2 Hz resulting in less tightly packed aggregates with reduced mechanical properties, probably due to the lack to kinetic energy required to overcome the initial friction between ribbons to rearrange after initial contact, Supplementary Figure S5.
@@
The shape of ribbons at a common temperature resulting from varying the offset angle is also expected to affect aggregate formation. We consider three different actuated shapes based on varying the offset angle, , exhibiting different amounts of curvature and twist (Supplementary Figure S2). As a control, we consider a polydomain ribbon, which does not exhibit any actuation (i.e., remains flat) at elevated temperature. Time lapse images of dispersions of ribbons with different shapes at 65∘C are shown in Figure 3d. Close-up images of aggregates resulting from ribbons with different offset angles that were entangled at 65∘C, are presented in Supplementary Figure S6b. The time evolution of the # of clusters for different offset angles at 65∘C is presented in Figure 3e, and the average number of ribbons per cluster is presented in Figure 3f. Again, the flapping motion performed by the polydomain ribbons (red line) did not result in any bending from the ribbons (Supplementary Video List of supporting videos) which did not result in entanglement.
@@
The kinetics of aggregation of different ribbon systems that aggregate did not exhibit any discernible differences. Although some variation is apparent, no substantial difference in kinetics or average cluster size was observed between the different angles at 65∘C. In all systems, two aggregation rates were observed, a fast drop in # of clusters lasting approximately 20 s followed by a slower decrease lasting approximately 80 s. In all systems, the formed aggregates occupied a volume approximately 80-90% smaller than in their dispersed state, Figure S7. This indicates that motion in circular patterns does result in successful entanglement. Similar observations were further reported by Patil et al. 13 in a computational study of active filaments representing dynamic reversible entanglement in the California Blackworms. We note that the worm bodies in the prior study frequently switch direction of rotation and that their bodies are highly flexible. By comparison, the ribbons follow a magnetic field rotating in only one direction and are somewhat rigid under the imposed forces. The evident considerable entanglement achieved in all systems in which the two of our preliminary conditions are fulfilled, i.e., continuous synchronicity with the field rotation and moderate curvature, can be considered to support our preliminary hypothesis.
Mathematical model for dynamic entanglement
@@
We use a mathematical model for the ribbon dynamics to identify the governing mechanisms of the aggregation process. In particular, this model allows us to understand the role of topological and non-topological mechanisms in aggregation. Our model treats the ribbons as rigid filaments (based on the observed change in curvature smaller than 10%) moving in a cylindrical domain with diameter and height , where is the length of a ribbon. To promote entanglement, the process is started from a random initial condition in which the center of mass of each ribbon is in the same plane. Each ribbon then experiences constant driving torques, analogous to the forces on the magnetic domain in the LCE ribbons, along with an external stochastic drift force. The ribbons also feel contact forces and torques arising from their interactions 32, 33 (Supplementary Information, Supplementary Video List of supporting videos). These contact forces typically include a cohesion term which prevents touching ribbons from separating unless their separation force exceeds a critical value. To understand how topological and non-topological parameters govern the aggregation process, we vary filament shape (Figure 4a; top row and bottom row) and cohesion strength (Figure 4a-c and Figure 4d-f).
@@
Filaments of different shapes exhibit different propensities to entangle with others, and thus help illustrate how topology affects aggregation. Here, two filament shapes are considered: a low curvature shape, corresponding to ribbons at C (Figure 4a, top row), and a high curvature shape, corresponding to ribbons at C (Figure 4a, bottom row). When cohesive forces are switched on, and the filaments are ‘sticky’, these different shapes give rise to strikingly different aggregation dynamics (Figure 4b) which resemble the experimental results in Figure 3b.
@@
The number of links in each filament provides a topological measure for successful entanglement. We introduce the excess linking per filament (Figure 4c), which quantifies the entanglements in a multi-filament system. This quantity is built out of the open linking number, , which captures pairwise entanglement between the ’th and ’th filament 13, 34. For filaments, the excess linking per filament is the average amount of linking with a single filament, relative to the average amount of linking found in a random configuration of filaments (Supplementary Information). At , the low curvature filaments exhibit a much more robust amount of entanglement than the high curvature filaments, as expected from their larger cluster sizes (Figure 4b). Previous work has indicated that an open linking number with magnitude approximately matches an intuitive notion of entanglement between two filaments 13. Since the high curvature filaments achieve an excess linking per filament of approximately (Figure 4c), this suggests that these filaments effectively pair off into clusters of size 2. This is consistent with the cluster number results (Figure 4b), which show the high curvature filaments forming clusters from initial filaments.